OpenStudy (marcelie):

help please find the area of the region . Decide whether to integrate with respect to x or y . Draw a typical approx rectangle and label the height and width.

OpenStudy (marcelie):

\[y= x^2-4x\] \[y=2x\]

OpenStudy (marcelie):

okay so i set them equal to each other in order to solve for x so how do i know if i would integrate from x or y ? \[x^2-4x-2x=0\] \[x^2-6x=0\] \[x=0 and x=6\]

OpenStudy (sshayer):

\[y=x^2-4x+4-4\] \[y+4=\left( x-2 \right)^2\] it is an upward parabola with vertex (2,-4)